OpenStudy (anonymous):

If you save three pennies on January 1, six pennies on January 2, nine pennies on January 3, and continue this pattern for one year (not a leap year), what will be the value of your entire savings, in dollars, at the end of that one year? Express your answer as a decimal.

OpenStudy (amistre64):

d:1 2 3 4 5 6 7 8 p:3 6 9 12 15 18 ...

OpenStudy (amistre64):

An = 3 + 3(n-1)

OpenStudy (amistre64):

3 + 3(364) = 3(1+364) 365 3 ----- 1095 we have a first terms of an arithmetic sequence of 3 and a last term at 1095, would you agree?

OpenStudy (amistre64):

\[Sum_n\frac{n(f+l)}{2}\]

OpenStudy (amistre64):

3,6,9,12,15; sums to 5(3+15)/2 = 5(9) = 45

OpenStudy (amistre64):

365(3+1095) ------------ = total sum 2

OpenStudy (anonymous):

Sn=(n/2)[2a+(n-1)d] total number of terms = 365 savings 3 6 9 12 ----------------- term(n) 1st 2nd 3rd 4th --------------- 365th n=no. of terms = 365 a=1st term = 3 d= common difference(between 2 terms) = 3 Sn=(365/2)[2*3+(365-1)3] =(365/2)[6+(364)3] =(365/2)[6+1092] =(365/2)(1098) =365*549 =200385

OpenStudy (amistre64):

should be 1095 .... 1*3 = 3 2*3 = 6 3*3 = 9 ..... 365*3 = 1095

OpenStudy (amistre64):

3 +...+1095 1095+...+ 3 ------------- 1098+...+1098, 1098(365)/2 looks like a read you wrong :)

OpenStudy (anonymous):

its not 3 3 3 3 3 ......(365 terms) but it is 3 6 9 12 -----(365) =3(1+2+3+4+5 -----+365)=3*66795 =200385

OpenStudy (amistre64):

you had it right, i just read something wrong in me head ... my fault :)

OpenStudy (anonymous):

2003.85